# SLMRec: Empowering Small Language Models for Sequential Recommendation

### v1

@@ Abstract
The sequential Recommendation (SR) task involves predicting the next item a user is likely to interact with, given their past interactions. The SR models examine the sequence of a user’s actions to discern more complex behavioral patterns and temporal dynamics. Recent research demonstrates the great impact of LLMs on sequential recommendation systems, either viewing sequential recommendation as language modeling or serving as the backbone for user representation. Although these methods deliver outstanding performance, there is scant evidence of the necessity of a large language model and how large the language model is needed, especially in the sequential recommendation scene. Meanwhile, due to the huge size of LLMs, it is inefficient and impractical to apply a LLM-based model in real-world platforms that often need to process billions of traffic logs daily. In this paper, we explore the influence of LLMs’ depth by conducting extensive experiments on large-scale industry datasets. Surprisingly, we discover that most intermediate layers of LLMs are redundant. Motivated by this insight, we empower small language models for SR, namely SLMRec, which adopt a simple yet effective knowledge distillation method. Moreover, SLMRec is orthogonal to other post-training efficiency techniques, such as quantization and pruning, so that they can be leveraged in combination. Comprehensive experimental results illustrate that the proposed SLMRec model attains the best performance using only 13% of the parameters found in LLM-based recommendation models, while simultaneously achieving up to 6.6x and 8.0x speedups in training and inference time costs, respectively.
@@ 1 Introduction
Learning temporal interest information is fundamental for sequential recommendation models. Traditional sequential recommendation (TSR) methods [52; 18; 29; 46] focus on the development of intricate sequential encoders, evolving from LSTM and GRU architectures to the self-attention layers and Transformer models. However, the state-of-the-art performance in TSR has hit a plateau, limited by model sizes that usually feature fewer than 0.1 billion parameters.
@@
Recently, Large Language Models (LLMs) [1; 47; 3] have made significant advancements in various aspects by scaling the size of the training data or the model’s architecture. Building upon the scaling laws delineated in prior research [30; 20], it endows Large Language Models (LLMs) with enhanced expressivity, culminating in superior performance benchmarks. Naturally, a burgeoning trend among contemporary LLM-based recommendation architectures has raised concerns. The current LLM-based recommender system can be classified as 1) generation-based approaches, e.g., P5 [12; 54], CoLLM [57] and LLaRa [38]; 2) embedding-based approaches such as E4SRec [35], CLLM4Rec [59] and Lite-LLM4Rec [50]. As shown in Fig. 1, generation-based approaches (G-LLMRec) encode an item as a token and formulate the sequential recommendation as the next token prediction task. By contrast, embedding-based approaches (E-LLMRec) regard the last hidden representation as user representation and learn an external adapter to compute user-item preference. The adoption of LLMs has vastly driven the development of sequence recommendation tasks, bringing an improvement of nearly 20% against the TSR model on the benchmark [35; 38; 50]. This arouses the following research motivation for this work.
$\bullet$ So far on the SR scene, little literature studies how large the language Model is needed. Recent studies [37; 14; 40] on the NLP domain suggest a high degree of redundancy in the LLM’s model architecture. Since the ID information has not been explicitly learned during the LLM’s training process, we also want to find out whether the scaling law of LLMs is also applicable to the SR task, specifically, how the relation is between the size of LLMs and their performance on SR task.
$\bullet$ Despite the large performance gain, the LLM-based SR solutions also escalate the model size significantly, e.g., nearly 70 times greater parameters compared with TSR models (from 0.1B to 7B+). Even within the parameter-efficient training technique [24], the paradigm still poses a significant challenge for real-world sequential recommendation use cases, where billions of traffic logs every day and potential new items need to be processed constantly. This disparity imposes strict hardware demands and makes it both inefficient and infeasible to deploy an LLM-based SR model.
@@
Figure 1: This overview compares traditional sequential recommendation (TSR) methods with LLM-based recommendation (LLMRec) methods. Here, $h_{u}$ and $h_{i}$ represent the user and item representations, respectively. In contrast to G-LLMRec methods [12; 54; 57; 38], E-LLMRec approaches [35; 59; 50] adhere to the TSR prediction framework. These methods leverage LLMs as feature extractors in the manner of BERT [8], diverging from the generative focus of G-LLMRec.
@@
Our contributions. This paper presents our initial attempt to reassess the need for large language models in sequential recommendation. To explore the reasons for the significant improvement, we conduct a series of experiments on large-scale industry datasets to investigate the effects of reducing the number of parameters during the training and inference stages on overall performance. From the empirical results, we found some profound insights that the improvement of the rise of the model parameters is not consistent. Meanwhile, it reveals that some layers of LLMs are redundant in the downstream recommendation task, similar to some recent findings in NLP domains [40; 14].
@@
Motivated by these findings, we empower small language models for the sequential recommendation, named SLMRec. We adopt the vanilla knowledge distillation approach [13; 6; 2] to align the representation knowledge. Moreover, multiple supervision signals are crafted to steer the student model toward acquiring task-aware knowledge within its hidden representations. Additionally, SLMRec operates without the need for any supplementary model design elements and is compatible with other quantization and pruning techniques utilized within Large Language Models (LLMs).
@@
Extensive experiments have revealed that SLMRec, with a model size of less than 1 billion parameters, can deliver performance that is remarkably competitive with baselines using LLMs sized over 7 billion parameters. Furthermore, SLMRec achieves up to 6.6x/8.0x speedup in terms of training/inference time costs against LLM-based recommendation models. Furthermore, we present the results of SLMRec employing online knowledge distillation, showing its competitive performance as well.
[…]
2 Motivation
@@
Evaluation Protocol. In the motivational experiment, we select SASRec as a traditional sequential recommendation baseline due to its performance [32]. We adopt [^1] To accelerate and follow the prediction head of traditional SR methods, we remove the original softmax layer and apply the dot product of user representation and item representation to obtain user-item prediction score. embedding-based method [35] as the baseline, named E4SRec, to easily generate the ranking for the full/sampled list of items. As shown in Fig. 2, a pre-trained embedding layer learned from SASRec is used to obtain the sequential item embedding. Then we concatenate the item embeddings with the prompt embeddings obtained after the tokenization. After encoding of stacked attention blocks of LLM, we regard the representation of the last layers as the user representation. Then, we follow the TSR methods to calculate the inner product of user embeddings and item embeddings from the pre-trained embedding layer to serve as the score for the user-item pair. Also, cross-entropy loss and fully candidate item are utilized for the optimization to achieve best results [53; 43]. To reduce both computational demands and processing time, LoRA [24] is used to update a comparatively smaller set of parameters. Besides, to generate an unbiased evaluation for fair comparison [33; 58], we randomly sampled 999 negative items, which were items not interacted with by the user, along with 1 positive item that served as the ground-truth interaction. To obtain large-scale industry data, we use the Amazon 18 version222 https://nijianmo.github.io/amazon/index.html dataset in this paper. More details are shown in Section 5.
[…]
@@ 3 Preliminaries
Figure 3: The overview of $\mathrm{SLMRec}$. A layer-wise knowledge distillation approach is applied to align the representation knowledge by grouping the layer into serveral blocks. The teacher and student model share a similar E-LLMRec model architecture. Multiple supervision signals are introduced to steer the student model toward acquiring fine-grained task-aware knowledge.
@@
Model structure. The E-LLMRec models capitalize on an ID embedding layer from TSR models such as BERT4Rec, SASRec, and GRU4Rec, which is pre-trained on a designated dataset [46; 29; 18]. The objective of sequential recommendation is to forecast subsequent items utilizing the user action sequence $S=(i_{1},i_{2},...,i_{T})$, a sequence that is either truncated or padded to maintain uniform length. Through truncation and padding, we derive the user’s action sequence mask, serving as the attention mask in LLMs (Large Language Models). The fixed-length sequence $S\in\mathbb{R}^{T}$ is translated into a sequential representation $\mathbf{S}\in\mathbb{R}^{T\times d_{0}}$ via the pre-trained ID embedding layer. A linear transformation is then applied to upscale the representation from a lower dimension $d_{0}$ to a higher dimension $d_{1}$ suitable for the hidden layers within the LLMs.
Upon defining the prompt template, the tokenization layer within the LLMs processes the natural language input into corresponding text embeddings and their associated attention masks. These embeddings and attention masks, derived from both the ID sequence and the text, are then introduced into the LLM decoder. The final temporal output $\mathbf{h}_{M}$ from the last layer of the decoder is inferred as the user representation and subsequently mapped through a linear layer to condense the dimensionality from $d_{1}$ back to $d_{0}$. Finally, user-item interaction predictions $\bar{p}$ are inferred by executing a dot product between the user and item representations. The learning process is refined through the application of a cross-entropy loss.
$\hat{p}_{i}=\frac{e^{\bar{p_{i}}}}{\sum\limits_{j\in I}e^{\bar{p_{j}}}};\;\;\;\mathcal{L}_{ce}=-\sum\limits_{u\in U,i\in I}y_{ui}log(\hat{p}_{i}).$ (1)
@@
Knowledge Distillation. Knowledge distillation is a technique aimed at transferring knowledge from a sophisticated teacher model to a more streamlined student model [19]. We represent the teacher by $f_{t}(\Theta^{t})$ and the student by $f_{s}(\Theta^{s})$. We aim to solve the following optimization problem:
$\min_{\Theta^{s}}[\mathcal{L}_{ce}(\Theta^{s})+\mathcal{D}_{kd}(\Theta^{t},\Theta^{s})].$ (2)
Here, $\mathcal{D}_{kd}(\Theta^{t},\Theta^{s})$ signifies the knowledge distillation loss, which quantifies the discrepancies between the teacher and the student models. A prevalent method involves employing the KL divergence to evaluate the divergence between the logits produced by both models. One well-established training schema is known as offline distillation, wherein the teacher is fully trained beforehand and remains unchanged, while the student is refined based on the criteria outlined in Eq. 6. With a designated training set, the teacher model $\Theta^{t}$ is initially trained by minimizing the cross-entropy loss $\mathcal{L}_{ce}$.
@@ 4 SLMRec
In this work, we do not conduct logits-based knowledge distillation since we want the student model to have the same ability to encode hidden representations rather than making similar predictions as the teacher. Therefore, we conduct feature distillation among every few layers. Considering the teach model has $M$ stacked decoder layers and student model has $N$ stacked decoder layers, we design several feature regularizars to conduct knowledge distillation between hidden representation from teach model $\mathbf{H}^{t}=\{\mathbf{h}^{t}_{m},...,\mathbf{h}^{t}_{m*i},...,\mathbf{h}^{t}_{b}\}$ and student model $\mathbf{H}^{s}=\{\mathbf{h}^{s}_{n},...,\mathbf{h}^{s}_{n*i},...,\mathbf{h}^{s}_{b}\}$ each few layers. Each $m$ layer of the teacher model and each $n$ layer of the student model are regarded as a block. The block number is $b$. In this work, we consider deeper LLaMa as the teacher model, and shallow LLaMa as the student model, which has the same hidden dimension between $\mathbf{H}^{t}\in\mathbb{R}^{b\times d_{1}}$ and $\mathbf{H}^{s}\in\mathbb{R}^{b\times d_{1}}$.
@@
$\mathcal{D}_{cos}(\Theta^{t},\Theta^{s})=\frac{1}{b}\sum\limits_{k\in b}\frac{\mathbf{h}^{t}_{k}\cdot\mathbf{h}^{s}_{k}}{\|\mathbf{h}^{t}_{k}\|_{2}\cdot\|\mathbf{h}^{s}_{k}\|_{2}}.$ (3)
@@
$\mathcal{D}_{norm}(\Theta^{t},\Theta^{s})=\frac{1}{b}\sum\limits_{k\in b}\|\mathbf{h}^{t}_{k}-\mathbf{h}^{s}_{k}\|_{2}^{2}.$ (4)
Multiple Supervision. Furthermore, we employ multiple supervision strategies to steer the student model toward assimilating specific aspects of recommendation-related knowledge. For each representation, we learn additional adapters ( $W_{t}$ ) to reduce the dimension. The modified prediction ( $\hat{p}^{mp}$ ) can be acquired as described by Eq. 1:
$\mathcal{L}_{ms}(\Theta^{s},W_{t})=\frac{1}{b}\sum\limits_{k\in b}\mathcal{L}_{ce}(y,\hat{p}^{mp}).$ (5)
@@
$\min_{\Theta^{s},W_{t}}[\mathcal{L}_{ce}(\Theta^{s})+\lambda_{1}\mathcal{D}_{cos}(\Theta^{t},\Theta^{s})+\lambda_{2}\mathcal{D}_{norm}(\Theta^{t},\Theta^{s})+\lambda_{3}\mathcal{L}_{ms}(\Theta^{s},W_{t})].$ (6)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are hyperparameters that govern the contribution of each term to the knowledge distillation process.

### v3

# SLMRec: Distilling Large Language Models into Small for Sequential Recommendation

@@ Abstract
Sequential Recommendation (SR) task involves predicting the next item a user is likely to interact with, given their past interactions. The SR models examine the sequence of a user’s actions to discern more complex behavioral patterns and temporal dynamics. Recent research demonstrates the great impact of LLMs on sequential recommendation systems, either viewing sequential recommendation as language modeling or serving as the backbone for user representation. Although these methods deliver outstanding performance, there is scant evidence of the necessity of a large language model and how large the language model is needed, especially in the sequential recommendation scene. Meanwhile, due to the huge size of LLMs, it is inefficient and impractical to apply a LLM-based model in real-world platforms that often need to process billions of traffic logs daily. In this paper, we explore the influence of LLMs’ depth by conducting extensive experiments on large-scale industry datasets. Surprisingly, our motivational experiments reveal that most intermediate layers of LLMs are redundant, indicating that pruning the remaining layers can still maintain strong performance. Motivated by this insight, we empower small language models for SR, namely SLMRec, which adopt a simple yet effective knowledge distillation method. Moreover, SLMRec is orthogonal to other post-training efficiency techniques, such as quantization and pruning, so that they can be leveraged in combination. Comprehensive experimental results illustrate that the proposed SLMRec model attains the best performance using only 13% of the parameters found in LLM-based recommendation models, while simultaneously achieving up to 6.6x and 8.0x speedups in training and inference time costs, respectively. Besides, we provide a theoretical justification for why small language models can perform comparably to large language models in SR. The source code and datasets are available at the URL  [^1] https://github.com/WujiangXu/SLMRec.
@@ 1 Introduction
Learning temporal interest information is fundamental for sequential recommendation models. Traditional sequential recommendation (TSR) methods (Wu et al., 2017; Hidasi et al., 2015; Kang & McAuley, 2018; Sun et al., 2019) focus on the development of intricate sequential encoders, evolving from LSTM and GRU architectures to the self-attention layers and Transformer models. However, the state-of-the-art performance in TSR has hit a plateau, limited by model sizes that usually feature fewer than 0.1 billion parameters.
@@
Recently, Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Anil et al., 2023) have made significant advancements in various aspects by scaling the size of the training data or the model’s architecture. Building upon the scaling laws delineated in prior research (Kaplan et al., 2020; Hoffmann et al., 2022), it endows LLMs with enhanced expressivity, culminating in superior performance benchmarks. Naturally, a burgeoning trend among contemporary LLM-based recommendation architectures has raised concerns. The current LLM-based recommender system can be classified as 1) generation-based approaches, e.g., P5 (Geng et al., 2022; Xu et al., 2023a), CoLLM (Zhang et al., 2023b) and LLaRa (Liao et al., 2023); 2) embedding-based approaches such as E4SRec (Li et al., 2023a), CLLM4Rec (Zhu et al., 2023) and Lite-LLM4Rec (Wang et al., 2024). As shown in Fig. 1, generation-based approaches (G-LLMRec) encode an item as a token and formulate the sequential recommendation as the next token prediction task. By contrast, embedding-based approaches (E-LLMRec) regard the last hidden representation as user representation and learn an external adapter to compute user-item preference. The adoption of LLMs has vastly driven the development of sequence recommendation tasks, bringing an improvement of nearly 20% against the TSR model on the benchmark (Li et al., 2023a; Liao et al., 2023; Wang et al., 2024). This arouses the following research motivation for this work.
@@
Figure 1: This overview compares traditional sequential recommendation (TSR) methods with LLM-based recommendation (LLMRec) methods. Here, $h_{u}$ and $h_{i}$ represent the user and item representations, respectively. In contrast to G-LLMRec methods, E-LLMRec approaches adhere to the TSR prediction framework. These methods leverage LLMs as feature extractors in the manner of BERT, diverging from the generative focus of G-LLMRec.
$\bullet$ Some researchers (Ardalani et al., 2022; Zhang et al., 2023a; 2024) have attempted to investigate the scaling laws in the recommendation domain. However, the largest model examined in these studies is less than 1 billion parameters, significantly smaller than the 175 billion parameters of GPT-3 (Brown et al., 2020). Additionally, the focus has been primarily on test loss rather than on ranking-based evaluation metrics, which limits the practical applicability of their findings. Recent studies (Liang et al., 2023; Gromov et al., 2024; Men et al., 2024) on the NLP domain suggest a high degree of redundancy in the LLMs’ model architecture. Since the ID information of the recommendation domain has not been explicitly learned during the LLMs’ training process, we also aim to find out whether increasing the model size of LLMs is beneficial for the SR task.
$\bullet$ Despite the large performance gain, the LLMRec methods also escalate the model size significantly, e.g., nearly 70 times greater parameters compared with TSR models (from 0.1B to 7B+). Even within the parameter-efficient training technique (Hu et al., 2021a), the paradigm still poses a significant challenge for real-world sequential recommendation use cases, where billions of traffic logs every day and potential new items need to be processed constantly. This disparity imposes strict hardware demands and makes it both inefficient and infeasible to deploy the LLMRec model.
@@
Our contributions. This paper presents an initial attempt to reassess the need for LLMs in sequential recommendation. To explore the reasons for the significant improvement of LLMRec methods, we conduct a series of experiments on large-scale industry datasets to investigate the effects of reducing the number of parameters during the training and inference stages on overall performance. From the empirical results, we found some profound insights that the improvement of the rise of the model parameters is not consistent. Meanwhile, it reveals that some layers of LLMs are redundant in the recommendation task, similar to findings in NLP domains (Men et al., 2024; Gromov et al., 2024).
@@
Motivated by these findings, we empower small language models for the sequential recommendation, named SLMRec. We adopt the vanilla knowledge distillation approaches to align the representation knowledge. Moreover, multiple supervision signals are crafted to steer the student model toward acquiring task-aware knowledge within its hidden representations. Additionally, our model operates without the need for any supplementary model design elements and is compatible with other quantization and pruning techniques utilized within LLMs.
@@
Extensive experiments have revealed that SLMRec, with a model size of less than 1 billion parameters, can deliver performance that is remarkably competitive with baselines using LLMs sized over 7 billion parameters. Furthermore, SLMRec achieves up to 6.6x/8.0x speedup in terms of training/inference time costs against LLM-based recommendation models. Besides, we present the results of SLMRec employing online knowledge distillation, demonstrating its competitive performance. Beyond empricial experiment results, we provide a theoretical justification for why small language models can perform comparably to large language models in SR.
(a) Cloth (Infer) - HR@10
(b) Cloth (Infer) - NDCG@10
(c) Cloth (Infer) - MRR
(d) Cloth (Train) - HR@10
(e) Cloth (Train) - NDCG@10
(f) Cloth (Train) - MRR
Figure 2: We present the relationship between the number of decoder layers and the final recommendation performance, with the performance of SASRec plotted as a baseline. Figures (a)-(c) show the results of directly using representations from the middle layers for inference without training, while (d)-(f) prune the later layers and train a model using only the specified number of layers. From the results, we observe that deeper decoder layers introduce redundancy in recommendation tasks, with models utilizing fewer layers (8-layer) achieving performance nearly equivalent to (24-layer) models.
2 Motivational Experiments
@@
Evaluation Protocol. In the motivational experiment, we select SASRec as a traditional sequential recommendation baseline due to its performance (Klenitskiy & Vasilev, 2023). We adopt [^2] To accelerate and align with the prediction head of traditional SR methods, we remove the original softmax layer and instead use the dot product of the user and item representations to compute the prediction score. embedding-based method (Li et al., 2023a) as the baseline, named E4SRec, to easily generate the ranking for the full/sampled list of items. As shown in Fig. 2, a pre-trained embedding layer learned from SASRec is used to obtain the sequential item embedding. Then we concatenate the item embeddings with the prompt embeddings obtained after the tokenization. After encoding of stacked attention blocks of LLM, we regard the representation of the last layers as the user representation. Then, we follow the TSR methods to calculate the inner product of user embeddings and item embeddings from the pre-trained embedding layer to serve as the score for the user-item pair. Also, cross-entropy loss and fully candidate item are utilized for the optimization to achieve best results (Xu et al., 2024a; Petrov & Macdonald, 2023). To reduce both computational demands and processing time, LoRA (Hu et al., 2021a) is used to update a comparatively smaller set of parameters. Besides, to generate an unbiased evaluation for fair comparison (Krichene & Rendle, 2020; Zhao et al., 2020), we randomly sampled 999 negative items, which were items not interacted with by the user, along with 1 positive item that served as the ground-truth interaction. To obtain large-scale industry data, we use the Amazon 18 version333 https://nijianmo.github.io/amazon/index.html dataset in this paper. More details are shown in Section 5.
Evaluation Strategy. To examine the connection between the number of parameters and the performance of LLM-based methods (E4SRec), we have truncated the original LLM architecture—in this case, a 32-layer decoder from the LLaMa 7B model—by pruning the decoder layers during both the inference and the training stages. As a direct inference method, we refrain from additional training using new labels and instead directly employ the output from the final ten layers as user representations to gauge recommendation performance. Instead of direct inference, we focus on conserving the initial layers of the decoder and proceed to train a more lightweight E4SRec model while adhering to the original training protocol. The models resulting from varying levels of layer retention are designated as E4SRecl, with the variable $l$ indicating the number of layers retained. The chosen values of $l$ encompass a spectrum, specifically $\{1,2,4,8,16,24,32\}$. Results from both experimental approaches are graphically depicted in Figure 2, providing insight into how the models’ depth influences their recommendation capabilities.
Insights. From Figure 2 (a)-(b), we can observe that directly utilizing the representation of other layers without training cannot obtain a comparative performance. Compared to TSR baseline SASRec, Figure 2 (c)-(d) yield the following insightful findings: (1) As the number of layers increases, the performance of the model also improves. Furthermore, even when the model has the same layer number (i.e., $l$=2) as SASRec, its performance is still superior to that of SASRec. We assume the gains observed in LLM-based methods could likely be attributed to the larger hidden representation size ((i.e., 4096 V.S. 128), the initialization from LLMs, and the introduction of PEFT (Hu et al., 2021a). (2) When $l$ is set ranging from 8-24, the model’s improvement is slight. It reveals that an 8-layer E4SRec8 can obtain nearly as informative user representations as a 24-layer E4SRec24. Considering the two findings above, it naturally inspires us to explore better training methods to obtain a smaller-size LLM-based SR model that is comparable with large models. If we want to learn a E4SRecM that perform similar as E4SRecN (M $<$ N), we should make sure the intermediate representations in E4SRecM to be as closer to those in E4SRecN as possible. Knowledge distillation (KD) is a straightforward idea in this case. Thus, we design a simple yet effective knowledge distillation method to train a tiny LLM-based model with similar performance. For the motivation experiment results in Movie domain, it can be found in Appendix B.1. In Section 6, we provide a theoretical justification that aligns with these empirical insights.
@@ 3 Preliminaries
Figure 3: The overview of SLMRec. A layer-wise knowledge distillation approach is applied to align the representation knowledge by grouping the layer into serveral blocks. The teacher and student model share a similar E-LLMRec model architecture. Multiple supervision signals are introduced to steer the student model toward acquiring fine-grained task-aware knowledge.
@@
Model structure. The E-LLMRec models capitalize on an ID embedding layer from TSR models such as BERT4Rec, SASRec, and GRU4Rec, which is pre-trained on a designated dataset (Sun et al., 2019; Kang & McAuley, 2018; Hidasi et al., 2015). The objective of sequential recommendation is to forecast subsequent items utilizing the user action sequence $S=(i_{1},i_{2},...,i_{T})$, a sequence that is either truncated or padded to maintain uniform length. Through truncation and padding, we derive the user’s action sequence mask, serving as the attention mask in LLMs (Large Language Models). The fixed-length sequence $S\in\mathbb{R}^{T}$ is translated into a sequential representation $\mathbf{S}\in\mathbb{R}^{T\times d_{0}}$ via the pre-trained ID embedding layer. A linear transformation is then applied to upscale the representation from a lower dimension $d_{0}$ to a higher dimension $d_{1}$ suitable for the hidden layers within the LLMs.
Upon defining the prompt template, the tokenization layer within the LLMs processes the natural language input into corresponding text embeddings and their associated attention masks. These embeddings and attention masks, derived from both the ID sequence and the text, are then introduced into the LLM decoder. The final temporal output $\mathbf{h}_{M}$ from the last layer of the decoder is inferred as the user representation and subsequently mapped through a linear layer to condense the dimensionality from $d_{1}$ back to $d_{0}$. Finally, user-item interaction predictions $\bar{p}$ are inferred by executing a dot product between the user and item representations. The learning process of the model is refined through the application of a cross-entropy loss.
$\hat{p}_{i}=\frac{e^{\bar{p_{i}}}}{\sum\limits_{j\in I}e^{\bar{p_{j}}}};\;\;\;\mathcal{L}_{ce}(\Theta_{s})=-\sum\limits_{u\in U,i\in I}y_{(ui)}log(\hat{p}_{i}).$ (1)
@@
Knowledge Distillation. Knowledge distillation is a technique aimed at transferring knowledge from a sophisticated teacher model to a more streamlined student model (Hinton et al., 2015). We represent the teacher by $f_{t}(\Theta_{t})$ and the student by $f_{s}(\Theta_{s})$. We aim to solve the following optimization problem:
$\min_{\Theta_{s}}[\mathcal{L}_{ce}(\Theta_{s})+\mathcal{D}_{kd}(\Theta_{t},\Theta_{s})].$ (2)
Here, $\mathcal{D}_{kd}(\Theta_{t},\Theta_{s})$ signifies the knowledge distillation loss, which quantifies the discrepancies between the teacher and the student models. A prevalent method involves employing the KL divergence to evaluate the divergence between the logits produced by both models. One well-established training schema is known as offline distillation, wherein the teacher is fully trained beforehand and remains unchanged, while the student is refined based on the criteria outlined in Eq. 6. In the offline knowledge distillation manner, the teacher model $\Theta_{t}$ is initially trained in a designated training set by minimizing the cross-entropy loss $\mathcal{L}_{ce}$.
@@ 4 SLMRec
In this work, we do not adopt logits-based knowledge distillation, as our goal is for the student model to learn how to encode hidden representations similar to the teacher model, rather than merely replicating its predictions. To achieve this, we perform feature distillation across multiple layers. Specifically, considering that the teacher model consists of $M$ stacked decoder layers and the student model has $N$ stacked decoder layers, we design several feature regularizers to guide the distillation process at regular intervals between the hidden representations of both models. We divide the layers of the teacher and student models into blocks by grouping every $m$ layers of the teacher and every $n$ layers of the student. The number of resulting blocks is $B$, calculated as $B=\left\lfloor\frac{M}{m}\right\rfloor=\left\lfloor\frac{N}{n}\right\rfloor$. Let the hidden representations from the teacher model be denoted as: $\mathbf{H}_{t}=\{\mathbf{h}_{t}^{m},\ldots,\mathbf{h}_{t}^{M}\}$, where $\mathbf{h}_{t}^{m}$ represents the final temporal dimension of the hidden representation from the $m$-th layer of the teacher. Similarly, the hidden representations from the student model are denoted as: $\mathbf{H}_{s}=\{\mathbf{h}_{s}^{n},\ldots,\mathbf{h}_{s}^{N}\}$. In this study, we use a deeper LLM as the teacher model and a shallower LLM as the student model, both sharing the same hidden dimension $d$, such that $\mathbf{H}_{t},\mathbf{H}_{s}\in\mathbb{R}^{B\times d}$.
@@
$\mathcal{D}_{\text{cos}}(\Theta_{t},\Theta_{s})=\frac{1}{B}\sum_{k=1}^{B}\frac{\mathbf{h}_{t}^{(km)}\cdot\mathbf{h}_{s}^{(kn)}}{\|\mathbf{h}_{t}^{(km)}\|_{2}\cdot\|\mathbf{h}_{s}^{(kn)}\|_{2}}.$ (3)
@@
$\vspace{-5pt}\mathcal{D}_{norm}(\Theta_{t},\Theta_{s})=\frac{1}{B}\sum_{k=1}^{B}\|\mathbf{h}_{t}^{(km)}-\mathbf{h}_{s}^{(kn)}\|_{2}^{2}.$ (4)
Multiple Supervision. Furthermore, we employ multiple supervision strategies to steer the student model toward assimilating specific aspects of recommendation-related knowledge. For each representation, we learn additional adapters ( $W_{a}$ ) to reduce the dimension. The modified prediction ( $\hat{p}_{t}^{(km)}$ ) can be acquired as described by Eq. 1:
$\vspace{-5pt}\mathcal{L}_{ms}(\Theta_{s},W_{a})=\frac{1}{B-1}\sum_{k=1}^{B-1}\mathcal{L}_{ce}(y,\hat{p}_{t}^{(km)}).$ (5)
@@
$\min_{\Theta_{s},W_{a}}[\mathcal{L}_{ce}(\Theta_{s})+\lambda_{1}\mathcal{D}_{cos}(\Theta_{t},\Theta_{s})+\lambda_{2}\mathcal{D}_{norm}(\Theta_{t},\Theta_{s})+\lambda_{3}\mathcal{L}_{ms}(\Theta_{s},W_{a})].$ (6)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are hyperparameters that control the contribution of each term.